Joseph is 50 inches tall and paul is y inches taller than jospeh and is 3 inches taller than dan. Write an expression for how much taller paul is than joseph
Answer by stanbon(75887)

Joseph is 50 inches tall and paul is y inches taller than jospeh and is 3 inches taller than dan. Write an expression for how much taller paul is than joseph
-----
j = 50
p = 50+y
p = d + 3
---
Equation:
p - j = 50+y - 50 = y inches